OpenStudy (anonymous):

Find the value of y so that the line passing through (2, 6) and (1, y) has a slope of 5

OpenStudy (anonymous):

use the slope formula: (y-6)/(1-2) = 5 (y-6)/(-1) = 5 multiply both sides by -1: y - 6 = -5 add 6 to both sides: y = 1

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

ur welcome!!!! :D please medal and fan! :)